# Noise Aware Utility Optimization of NISQ Devices

### v1

@@ II Background & Related Work
NISQ-era devices are highly affected by noise [2][3]. Moreover, it has been shown by [4] that each IBM quantum computer exhibits a time dependant noise fingerprint so distinct it can be learnt. As depicted in Figure 1, there is a discernible increase in the average CNOT error rate correlated with the age of a sample of devices from IBM’s fleet. One contributing factor to this trend is the drift in the resistance [5] of the Josephson Junctions used to create qubits as they age, resulting in qubit frequency fluctuations. We believe those fluctuations are responsible, in part, for the degradatation of gate fidelity on aging devices. Additional studies, including by [6] and [7], independently confirm that T1 relaxation time variability of qubits is a primary contributor to the frequency and dephasing fluctuations, as well as to the observed variations in gate fidelity. These findings underscore the importance of addressing time-dependent qubit performance degradation for achieving reliable quantum computations.
@@ IV-C Experimental Setup
Our experimental setup consists of generating random paths across the largest partitions obtained post-pruning, employing a random walk algorithm. Each of those paths will later lead to one circuit being executed on the system. By invoking the initial_layout flag, we direct Qiskit’s transpiler to conform its transpilation process to the generated partitions, ensuring results integrity. Quantum Process Tomography is then used to estimate the overall process fidelity. A simple linear relationship is used to convert those results to gate fidelity: $F_{\textrm{gate}}=(4F_{\textrm{process}}+1)/5$ [14][15]. Finally, the experiments were run over the course of three weeks on ibm_quebec, totalling 3,000+ circuit executions of 4096 shots each. Overall, those experiments took about 492,000 Qiskit Runtime seconds–roughly 135 hours.

### v2

@@ II Background & Related Work
NISQ-era devices are highly affected by noise [2][3]. Furthermore, research by [4] demonstrates that each IBM quantum computer has a unique, time-dependent noise signature that can be identified and learned over time. As depicted in Figure 1, there is a discernible increase in the average CNOT error rate correlated with the age of a sample of devices from IBM’s fleet. One contributing factor to this trend is the drift in the resistance [5] of the Josephson Junctions used to create qubits as they age, resulting in qubit frequency fluctuations. We believe those fluctuations are responsible, in part, for the degradatation of gate fidelity on aging devices. Additional studies, including by [6] and [7], independently confirm that T1 relaxation time variability of qubits is a primary contributor to the frequency and dephasing fluctuations, as well as to the observed variations in gate fidelity. These findings underscore the importance of addressing time-dependent qubit performance degradation for achieving reliable quantum computations.
@@ IV-C Experimental Setup
Our experimental setup consists of generating random paths across the largest partitions obtained post-pruning, employing a random walk algorithm. Each of those paths will later lead to one circuit being executed on the system. By invoking the initial_layout flag, we direct Qiskit’s transpiler to conform its transpilation process to the generated partitions, ensuring results integrity. Quantum Process Tomography (QPT) is used to estimate the fidelity of quantum gates by reconstructing the quantum process from measurements. This reconstruction involves comparing the theoretical and experimental process matrices, providing insight into how closely the quantum device performs to expectations. Once an estimate of the process fidelity is obtained, a simple linear relationship is used to convert it to gate fidelity: $F_{\textrm{gate}}=(4F_{\textrm{process}}+1)/5$ [14][15]. Gate fidelity is chosen because it provides a scalar value that encapsulates the overall performance of the quantum gate, making it a relevant figure of merit for our method’s effectiveness. Finally, the experiments were run over the course of three weeks on ibm_quebec, totalling 3,000+ circuit executions of 4096 shots each. Overall, those experiments took about 492,000 Qiskit Runtime seconds–roughly 135 hours.